### v4

@@ 1 Introduction
Low-polynomial variants.  Only a few Art Gallery variants are solvable in low-polynomial time, and all known results rely on very strong structural restrictions. Lee and Preparata [Lee1979ptimal] gave a linear-time algorithm to decide whether a polygon can be guarded by a single guard. De Berg, Durocher, and Mehrabi. [DBLP:journals/jda/BergDM17] presented a linear-time algorithm to minimize the number of edge guards under orthogonal vision when $P$ an rectilinear polygon. Palios and Tzimas [Palios2013-star-cover] restrict each guard’s visibility region to an orthogonal $r$-star and restrict the polygon to class-3 orthogonal polygons, and obtained near-linear time. Biedl and Mehrabi [BiedlMehrabi2016_rGuardingThinOrthogonal] combined several restrictions—$P$ must be orthogonal [SchuchardtHecker1995], thin [Tomas2013ThinOrth], and vision rectilinear [Biedl2017Mehrabi].
@@ 5 Intermezzo: data structures and sliding sequences
  • The path $S(s,t)$ is represented by $O(\log n)$ balanced trees whose ordered leafs traverse $S(s,t)$.
@@ 6 Computing all reduced dominators
Consider the polygon $P^{\prime}$ bounded by $S_{s}$, $[s,t]$, and $S_{t}$. Since $g$ sees $m$, the segment $\overline{g\,m}$ splits $P^{\prime}$ into two polygons, $P_{s}\subseteq P$ and $P_{t}\subseteq P$, containing $s$ and $t$ respectively (see Figure˜9). Let $u^{*}$ be the maximal point in $[s,m]$ such that $\overline{g\,u^{*}}$ intersects $S_{s}$ in more than one point (i.e., not only in $g$). We first show that $g$ sees all of $u^{*},m]$. Indeed, consider the polygon $P_{u^{*}}$ bounded by $u^{*},m]$, $\overline{m\,g}$, and $\overline{g\,u^{*}}$. Because $g\in\mathcal{E}[s,t]$, it lies left of all edges of $P_{u^{*}}$ and thus, by Section˜2, $g$ sees all of $u^{*},m]$. By construction, $P_{u^{*}}$ contains no point of $S_{s}$ in its interior and hence lies within $P_{s}\subseteq P$. Therefore, $g$ indeed sees $u^{*},m]$ in $P$.
@@ 9 An $O(n\log n)$-time algorithm
Given the presented results, we can finally present our main result. After $O(n\log n)$ time we can either output a solution of size at most $3$, or have a representation of next as a piecewise Möbius transform, consisting of $O(n)$ pieces. Note that given this data, we can get a $(+1)$-approximation $k^{\prime}$ of $k$ in $O(n)$ time by repeatedly applying $\textnormal{{next}}(\cdot)$ to an arbitrary $x\in\partial P$ [RobsonSpaldingZheng2024_AnalyticArcCover, Lemma 7] via a linear sweep of the representation of next. We finally use the techniques of [aggarwal1989finding, Section 4], which in $O(n\log k)$ checks if there is any point $x\in\partial P$ such that $\textnormal{{next}}^{k^{\prime}-1}(x)\geq x+n$.

### v5

@@ 1 Introduction
Low-polynomial variants.  Only a few Art Gallery variants are solvable in low-polynomial time, and all known results rely on very strong structural restrictions. Lee and Preparata [Lee1979ptimal] gave a linear-time algorithm to decide whether a polygon can be guarded by a single guard. De Berg, Durocher, and Mehrabi [DBLP:journals/jda/BergDM17] presented a linear-time algorithm to minimize the number of edge guards under orthogonal vision in a a monotone rectilinear polygon. Palios and Tzimas [Palios2013-star-cover] restrict each guard’s visibility region to an orthogonal $r$-star and restrict the polygon to class-3 orthogonal polygons, and obtained near-linear time. Biedl and Mehrabi [BiedlMehrabi2016_rGuardingThinOrthogonal] combined several restrictions—$P$ must be orthogonal [SchuchardtHecker1995], thin [Tomas2013ThinOrth], and vision rectilinear [Biedl2017Mehrabi].
@@ 5 Intermezzo: data structures and sliding sequences
  • The path $S(s,t)$ is represented by $O(\log n)$ balanced trees whose ordered leaves traverse $S(s,t)$.
@@ 6 Computing all reduced dominators
Consider the polygon $P^{\prime}$ bounded by $S_{s}$, $[s,t]$, and $S_{t}$. Since $g$ sees $m$, the segment $\overline{g\,m}$ splits $P^{\prime}$ into two polygons, $P_{s}\subseteq P$ and $P_{t}\subseteq P$, containing $s$ and $t$ respectively (see Figure˜9). Let $u^{*}$ be the maximal point in $[s,m]$ such that $\overline{g\,u^{*}}$ intersects $S_{s}$ in more than one point (i.e., not only in $g$). We first show that $g$ sees all of $u^{*},m]$. Indeed, consider the polygon $P_{u^{*}}$ bounded by $u^{*},m]$, $\overline{m\,g}$, and $\overline{g\,u^{*}}$. Because $g\in\mathcal{E}[s,t]$, it lies left of all edges of $P_{u^{*}}$ and thus, by Section˜2, the visibility of $g$ to $u^{*},m]$ is not blocked by any edge in $[s,t]$. . By construction, $P_{u^{*}}$ contains no point of $S_{s}$ in its interior and hence lies within $P_{s}\subseteq P$. Therefore, $g$ indeed sees $u^{*},m]$ in $P$.
@@ 9 An $O(n\log n)$-time algorithm
Given the presented results, we can finally present our main result. After $O(n\log n)$ time we can either output a solution of size at most $3$, or have a representation of next as a piecewise Möbius transform, consisting of $O(n)$ pieces. Note that given this data, we can get a $(+1)$-approximation $k^{\prime}$ of $k$ in $O(n)$ time by repeatedly applying $\textnormal{{next}}(\cdot)$ to an arbitrary $x\in\partial P$ [RobsonSpaldingZheng2024_AnalyticArcCover, Lemma 7] via a linear sweep of the representation of next. We finally use the techniques of [aggarwal1989finding, Section 4], which in $O(n\log k)$ checks if there is any point $x\in\partial P$ such that $\textnormal{{next}}^{k^{\prime}-1}(x)\geq x+n$. We wish to give maximal credit to [aggarwal1989finding, Section 4] as it is able to directly resolve this subproblem. However, as a reviewer has questions the validity of this application, we note that it is straightforward to find an $O(n\log n)$-time algorithm that is derived from first principles and we present this in Section A
@@ 11 Conclusion
Appendix A Deriving the function evaluation from first principles
Consider traversing the polygon boundary twice, and denote the traversal as the interval $[1,2n+1]$ (where each integer corresponds to a vertex of the polygon and each vertex appears twice). By Lemma 6.7, we obtain in $O(n\log n)$ time a linear-size set of points $X\subset[1,n+1)$ such that there exists an optimal solution (of size $k$) that is applied by applying the next function $k$ times to some $x\in X$. What remains, is to compute the minimum integer $k$ such that there exists an $x\in X$ where applying the next function $k$ times “overtakes” $x$ (i.e., $\textnormal{{next}}^{k}(x)\geq n+x$.
Our analysis in this paper shows that we can partition $[1,2n+1]$ into intervals $I_{1},\ldots,I_{m}$ with $m\in O(1)$ such that for all $x\in I_{i}$ there exists a constant-complexity Möbius transform $f_{i}(x)$ that maps $x$ to $\textnormal{{next}}(x)$. We have showed that, by computing all good dominators, bad dominators, and all functions representing ugly dominators, that $m\in O(n)$ and that we can compute the pairs $(I_{i},f_{i})$ in $O(n\log n)$ time. Given the set of functions $\{(I_{i},f_{i})\}$ and the set of $O(n)$ starting points $X\subset[1,n+1)$, we wish to evaluate these functions on $X$. These functions have two crucial properties:
  • For any two functions $(f_{i},f_{j})$ where $f_{i}(x)\in I_{j}$, the compound $f_{i}\circ f_{j}$ is another constant-complexity function that, on a Real RAM, can be evaluated in constant time. Indeed, if the guard corresponding to $f_{i}$ and $f_{j}$ are ugly dominators then the compound of two Möbius transforms is another Möbius transform and if $f_{j}$ is a good or bad dominator then the resulting is a single number $v$ where $v$ is the right endpoint of the corresponding guard $(g,[u,v])$.
  • The points in $X$ stay in-order when recursively applying the next function. In particular, consider some $x\in X$ with $x\in X_{i}$ and suppose that, at some point, applying next projects $x$ onto $I_{j}$ in between points $y\in I_{j}\cap X$ and $z\in I_{j}\cap X$. Then as we recursively apply the next to $y$, $z$, and the projection of $x$ these points will stay in the order that has the projection of $y$ first, followed by the projection of $x$, followed by the projection of $z$.
These two properties allow us to evaluate next recursively by using lazy evaluation along a segment tree that stores $X$. In particular, partition $X$ such that $X_{i}$ denotes the subset of $X$ that is contained in $I_{i}$. We iterate over integers $i\in[1,m]$ (note that this means we go around the polygon twice). We maintain an “evaluation tree” $T$. It stores in its leaves all the points in $X_{j}$ for $j<i$. We then define the evaluated point $F_{i}(x)$, which is the first point obtained by applying next recursively from $x$ until it exceeds the left boundary of $I_{i}$. We maintain the following three invariants at each iteration $i$:
  1. 1.
    In the inner nodes of the tree, we store a single function called a symbol. A symbol is either the identify function, or the compound of functions $f_{j}$ for $j<i$. We maintain the invariant that for each leaf $x$ of $T$, the compound of all symbols on its leaf-to-root path, applied to $x$, yields $F_{i}(x)$.
  2. 2.
    We also maintain the invariant that the leaves are sorted by $F_{i}(x)$. That is, for a leaf $x$ left of a leaf $y$, $F_{i}(x)$ lies before $F_{i}(y)$.
  3. 3.
    Finally, we (implicitly) maintain for each leaf $x\in T$ the interval $I_{j}$ where $j\geq i$ that contains $F_{i}(x)$ and, for this interval, we explicitly maintain the first and last leaf $y$ and $z$ in this interval plus their evaluated points $F_{i}(y)$ and $F_{i}(z)$. We obtain this by marking for each $I_{j}$ the contiguous subtrees of $T$ that contain all leaves $x\in T$ where $F_{i}(x)\in I_{j}$.
These three invariants are maintained precisely by the two highlighted properties of our functions we mentioned before. In particular, our algorithm iterates over integers $i\in[1,m]$ and maintains these invariants in three steps. The first step checks if $I_{i}$ intersects $[1,n+1]$ (and we have thus not finished the first round around the polygon yet) and, if so, inserts all points of $X_{i}$ into $T$ and restores our invariant. The second step then applies the function $f_{i}$ to all leaves $x\in X$ for which $F_{i}(x)$ lies in $I_{i}$ to ensure that the invariant holds at the start of iteration $i+1$. The third step ensures that we update the markings in our subtree. I.e., that we correctly record for all $I_{j}$ the contiguous subtrees of $T$ that contain all leaves $x\in T$ where $F_{i}(x)\in I_{j}$. Figure 22 illustrates our approach.
The first step.
Because the leaves of $T$ are ordered according to our second invariant, it must be that all the leaves in $T$ for which $F_{i}(x)$ lies in the interval $I_{i}$ form a contiguous set $X^{\prime}$. Our third invariant identifies $X^{\prime}$ as $O(\log n)$ subtrees of $T$. We use the identification of $I_{i}$ to insert each point $x\in X_{i}$ into $T$ as a leaf, where $x$ is placed in-between the two consecutive leaves $a,b\in T$ where $F_{i}(a)$ precedes $x$ and $F_{i}(b)$ succeeds $x$ (we can achieve this through invariant 3 by binary searching over all marked intervals $I_{j}$ in the tree, by comparing $x$ to the stored points $F_{i}(y)$ and $F_{i}(z)$. This finds us the subtrees that correspond to the interval $I_{i}$. From there, we note that we can insert $N$ values in $O(N\log n)$ time by comparing to the root-to-leaf paths of the leaves in these subtrees directly). This insertion does not break the second and third invariant. To restore our first invariant, we must ensure that for the new leaf $x$ its root-to-leaf path contains no symbols. We use a “push down” procedure on the symbols we encounter during a root-to-leaf traversal to $x$: traverse the root to $x$, one node at a time. At each node $v$, compound the symbol at $v$ into the symbols seen so far. If $x$ lies in the left child of $v$, compound the result into the right child of $v$ and erase the symbol at $v$ (maintaining the first invariant for that subtree). If $x$ lies in the right child of $v$, we naturally compound the result into the left child instead. This procedure inserts a point $x\in X_{i}$ into $T$ in $O(\log n)$ time. Since each point is inserted exactly once, this terminates in $O(n\log n)$ time.
The second step.
Now that all points in $X_{i}$ are inserted into $T$, there is a contiguous interval $X^{\prime}$ in $T$ that contains all the points whose evaluated point $F_{i}$ lies in $I_{i}$. We want to (implicitly) apply $f_{i}$ to all these points. As these leaves from a contiguous interval, there are $O(\log n)$ nodes in our tree that together represent $X^{\prime}$. We want to add the symbol for $f_{i}$ to these $O(\log n)$ nodes. However, it is crucial that our symbols stay in-order. That is, in a leaf-to-root path for any element in $X^{\prime}$ must have $f_{i}$ appear at its last symbol. We observe that if we take the union $U$ of all paths from the root to a node $v$ that we want to add $f_{i}$ to, then the total number of tree nodes in $U$ is $O(\log n)$. To maintain this order, we again use a “push down” strategy across nodes in $U$. That is, if we wish to add the symbol for $f_{i}$ to a node $v^{*}$ then we traverse the root to $v^{*}$, making sure that we never traverse a node in $U$ twice. If we encounter a node $v\neq v^{*}$ then we read the symbol at that node. We compound the node into its two children and erase the symbol at $v$. This maintains our invariant until we reach $v^{*}$, where at $v^{*}$ there now is one symbol that is the compound of the path from $v^{*}$ to the root. We compound $f_{i}$ with this symbol to maintain our first invariant for $i+1$. Note that the second invariant is not affected by this procedure. Since there are $O(\log n)$ nodes across the union of root-to-leaf paths that receive a new symbol, we can charge the $O(\log n)$ time spent to the interval $I_{i}$. Since each interval $I_{i}$ is charged at most twice, this requires $O(n\log n)$ total time.
The third step.
What remains is to maintain our third invariant and we show that our third invariant also suffices to test for the first time when a point overtakes itself. Whenever we apply the function $f_{i}$, the points under this function get mapped to intervals $I_{j_{1}},I_{j_{2}},\ldots$ for integers $j$ bigger than $i$. We say that the points in $X^{\prime}$ gets split across these intervals. Denote by $I_{j^{*}}$ the last interval and observe that for all future iterations $i^{\prime}$, all points get mapped to $I_{j^{*}}$ or further. This means that for all intervals in the split, except $I_{j^{*}}$, we can charge $O(\log n)$ time to find out the contiguous subset of $X^{\prime}$ that gets mapped to $I_{j^{*}}$ and we update the marking in $T$. Specifically, we can consider any node whose subtree contains only leaves in $X^{\prime}$ and we can evaluate $F_{i}(a)$ and $F_{i}(b)$ for the leftmost and rightmost leaf under that node in $O(\log n)$ time to detect whether this node “splits”. If it does not, then we terminate and charge $O(\log n)$ time to the unique interval $I_{j}$ that all these points land into. If it does, we charge to the split and there can only be as many splits as there are intervals. As each marked interval $I_{j}$ contains only one leftmost and one rightmost leaf $y$ and $z$, we can evaluate $F_{i}(y)$ and $F_{i}(z)$ in $O(\log n)$ time each by traversing their leaf-to-root paths, charging this time spent to $I_{j}$. We also update the marking for $I_{j^{*}}$ in the same manner, charging it to the iteration $i$ instead.
Given that we know for each contiguous set of leaves which new interval $I_{j}$ it lands in, we can determine in $O(\log n)$ time per new interval whether there exists a leaf that lands in $I_{j}$ whose start point +n+n is lower than the left endpoint of $I_{j}$ in exactly the same way as interval trees can get a minimum value in an interval in logarithmic time.
Figure 22: The figure illustrating the algorithm used to maintain our evaluation tree.